HUBBLE SPACE TELESCOPE DAILY REPORT #2146 PERIOD COVERED: 0000Z (UTC) 06/16/98 - 0000Z (UTC) 06/17/98 Daily Status Report as of 168/0000Z 1.0 ACTIVITIES SCHEDULED AND ACCOMPLISHED: 1.1 Completed STIS/CCD 7926 (CCD Dark and Bias Monitor -- Continued) The Space Telescope Imaging Spectrograph (CCD) was used to make several dark and bias calibration observations. This is to measure CCD dark current and bias rates for the two supported amplifier settings {gain = 1 and gain = 4} in order to gather data for making superdarks and superbiases. These observations will also track the growth of hot pixels on the CCD in general and within the various target acquisition apertures specifically. During this proposal, losses of lock (enumerated in HSTAR 6542) occurred at 167/001615Z and 167/002856Z. The observations were executed as scheduled, and no further anomalies were noted. 1.2 Completed STIS/CCD 7912 (STIS Parallel Archive Proposal - Nearby Galaxies - Imaging and Spectroscopy) The Space Telescope Imaging Spectrograph (CCD) was used to make parallel observations of nearby galaxies. This survey will be useful to study the star formation histories, chemical evolution, and distances to these galaxies. These data will be placed immediately into the Hubble Data Archive. During this proposal, a loss of lock (enumerated in HSTAR 6542) occurred at 167/015154Z. The observations were completed as scheduled, and no further problems were reported. 1.3 Completed STIS/CCD 7908 (POMS Test Proposal: STIS Non-Scripted Parallel Proposal) The Space Telescope Imaging Spectrograph (CCD) was used to make calibration observations as part of the POMS Test Proposal. This was a STIS non-scripted parallel proposal. During this proposal, losses of lock (enumerated in HSTAR 6542) occurred at 167/020538Z and 167/021819Z. The observations were completed as planned, and no problems were reported. 1.4 Completed STIS/MA2 7604 (Cycle 7 MAMA Dark Measurements) The Space Telescope Imaging Spectrograph (MA2) was used to make MAMA dark measurements. These darks will be used in later image reductions and also trended over the instrument's life time as an indicator of the detector's degradation or damage. The observations were completed as scheduled, and no problems were reported. 1.5 Completed NIC/3 7919 (NICMOS Camera 3 Campaign Snapshots) The Near Infrared Camera (Camera Three) was used to make images of galaxies NGC-2841 and NGC-3077 as part of the NICMOS Camera 3 campaign snapshots. The proposal is to obtain wide field snapshot images of nearby galaxies in Paschen Alpha which is inaccessible from the ground. The observations were executed as planned, and no anomalies were reported. 1.6 Completed Two Sets of WF/PC-2 7909 (POMS Test Proposal: WF II Parallel Archive Proposal) The WF/PC-2 was used to make parallel observations as part of the POMS test proposal. This test proposal is designed to simulate scientific plans. The observations were completed as planned, and no problems were reported. 1.7 Completed Two Sets of STIS/CCD 7635 (STIS CCD Hot Pixel Annealing) The Space Telescope Imaging Spectrograph (CCD) was used to test the hot pixel annealing procedure. The effectiveness of the CCD hot pixel annealing process is assessed by measuring the dark current behavior before and after annealing and by searching for any window contamination effects. The observations were completed as planned, and no problems were reported. 1.8 Completed Two Sets of NIC/1/2/3 7817 (A Complete NICMOS Map of the Hubble Deep Field) The Near Infrared Camera (Cameras One, Two and Three) was used to make observations to map the complete WF/PC-2 region of the central Hubble Deep Field. These data will provide unique information about the faintest objects, but will include only a small number of objects with spectroscopic redshifts or z ~ 3 Lyman break galaxies. During the first iteration of this proposal, losses of lock (enumerated in HSTAR 6542) occurred at 167/072755Z, 167/073139Z, and 167/090441Z. The observations were executed as scheduled, and no further anomalies were noted. 1.9 Completed Two Sets of WF/PC-2 8013 (WF/PC-2 Parallel Images to the HDF Flanking Fields {Companion Proposal to 7817}) The WF/PC-2 was used to make observations parallel to the Hubble Deep Field observations (i.e., flanking fields). The complete WF/PC-2 region of the central Hubble Deep Field {HDF} was being mapped with the NICMOS Camera 3 at F110W and F160W {proposal 7817}. While these observations were executing, the WF/PC-2 was working in parallel mode to image areas near this field. During the first iteration of this proposal, losses of lock (enumerated in HSTAR 6542) occurred at 167/072755Z, 167/073139Z, and 167/090441Z. The observations were completed as scheduled, and no further anomalies were reported. 1.10 Completed S/C 5582 (FOC/48 Monthly Activation (Camera Section Only): Cycle 4) This proposal was executed as scheduled, and no anomalies were noted. 1.11 Completed NIC/1/2/3 7322 (The IMF Below 1 Msun in Young Clusters: 1.9 Micron Water Band Imaging of IC348) The Near Infrared Camera (Cameras One, Two, and Three) was used to make observations of the cluster around HD196795. This study will be the first to obtain spectral types for a young cluster population complete to K=16, and thereby quantify, to unprecedentedly high precision, the IMF below 1M{sun} in young clusters. The observations were executed as scheduled, and no anomalies were reported. 1.12 Completed NIC/3 7892 (Narrow-band Imaging of BAL QSOs) The Near Infrared Camera (Camera Three) was used to make narrow band observations of the BL-Lac (BAL) PHL5200. The detection of extended emission-line gas structures in BAL QSOs would offer another means of investigating the BAL QSO phenomenon as well as establishing a new probe of material in galaxies at high redshift. The observations were completed as planned, and no problems were reported. 1.13 Completed NIC'/37858 (The Composition of Kuiper Belt Objects) The Near Infrared Camera (Camera Three) was used to investigate the small bodies recently discovered in the outer solar system (i.e., A1996TP66) that appear to be unique both in composition and diversity compared to other classes of solar system objects. The proposal completed as planned, and no anomalies were noted. 2.0 FLIGHT OPERATIONS SUMMARY: 2.1 Guide Star Acquisitions: Scheduled Acquisitions :9 Successful :9 Scheduled Reacquisitions : 9 Successful : 9 Losses of lock reported in yesterday's HSTAR 6542 affected proposals described in 1.2, 1.3, 1.8, and 1.9 above. HSTAR 6544 reported a loss of lock during a type 3 slew at 168/0916Z. NICMOS proposal 7817 (A Complete NICMOS Map of the Hubble Deep Field) and WF/PC-2 proposal 8013 (WF/PC-2 Parallel Images to the HDF Flanking Fields), both occurring during the period of tomorrow's report, were affected by this event. 2.2 FHST Updates: Scheduled : 10 Successful : 10 2.3 Operations Notes: Using ROP SR-1, the SSR EDAC error counter was cleared on two occasions. As directed by operation note #444, the K-based recharge ratio for Sun orbit time was adjusted at 167/2115Z. The MAMA-1 threshold limit was changed at 168/0400Z, using operations note #443 Per an operations request, a real-time FHST map using trackers one and two was performed at 168/0556Z. 3.0 SIGNIFICANT FORTHCOMING EVENTS: Continuation of normal science observations and calibrations. /DMH